Simplify.
1/3x - 4/3= -1/6x - 1

Answer by unlockmath(1688)

Hello,
Let's first get rid of the fractions by multiplying it by 6x. (1/3x - 4/3= -1/6x - 1)
This will then be:
2-8x=-1
Now subtract 2 from both sides to get:
-8x=-3
Now divide by -8 to get:
x=3/8 (Hint: Simplify deals with expressions. Solving has to do with equations)
There we go. Make sense?
RJ
www.math-unlock.com